Sara's river boat travels downstream 16 miles in 2 hours. It returns the same distance upstream in 3 hours. Find the rate of the boat in still water and the rate of the current. Please explain how to get the answer. Thank you
Answer by venugopalramana(3286)

SEE THIS WHICH IS SIMILAR TO YOUR PROBLEM AND SOLVE
YOUR PROBLEM...NOTE THAT STILL WATER MEANS VELOCITY OF STREAM=0
A boat travels at the speed of 20 Miles per hour in
still water. It travels 48 miles up stream, and then
returns to the starting point in a total of five
hours. The speed of the current is ____ miles per
hour.
distance =speed * time..or...time=distance/speed
Still water speed of boat =20 mph
let the stream velocity be =x mph
so its speed upstream =20-x
its speed down stream =20+x
distance travelled either way =48 miles
time taken to travel upstream=48/(20-x)
time taken to travel down stream =48/(20+x)
total time of travel =5 hours..hence
[48/(20-x)]+[48/(20+x)]=5
48[(20+x+20-x)/((20-x)(20+x))]=5
48*40=5(20^2-x^2)=2000-5x^2
5x^2=2000-1920=8=
x^2=80/5=16